### 29 Nov 2023 v1

# Quantum Neural Networks under Depolarization Noise: Exploring White-Box Attacks and Defenses

Abstract
Leveraging the unique properties of quantum mechanics, Quantum Machine Learning (QML) promises computational breakthroughs and enriched perspectives where traditional systems reach their boundaries. However, similarly to classical machine learning, QML is not immune to adversarial attacks. Quantum adversarial machine learning has become instrumental in highlighting the weak points of QML models when faced with adversarial crafted feature vectors. Diving deep into this domain, our exploration shines light on the interplay between depolarization noise and adversarial robustness. While previous results enhanced robustness from adversarial threats through depolarization noise, our findings paint a different picture. Interestingly, adding depolarization noise discontinued the effect of providing further robustness for a multi-class classification scenario. Consolidating our findings, we conducted experiments with a multi-class classifier adversarially trained on gate-based quantum simulators, further elucidating this unexpected behavior.
Keywords Quantum Machine Learning  $\cdot$ Quantum Computing  $\cdot$ Adversarial Robustness
1 Introduction
The field of quantum computing has shown fast advancements in recent years, providing the possibility to speed up speficic tasks. Hereby a variety of algorithms exists such as Shor’s algorithm to factor primes [30], Grover’s search to find an element in an unsorted list in $\mathcal{O}(\sqrt{n})$ [10] or the HHL-Algorithm to efficiently solve systems of linear equations [11]. In this context, quantum machine learning (QML) focuses onto providing the properties and speedups of quantum computing in the field of machine learning [33, 32, 28, 1]. First coined by Kak [13], the term quantum neural network (QNN) emerged as an ambitious effort to merge observations from the domain of neuroscience with the intriguing features inherent to quantum computations. Essentially, QNNs represent a specific type of Quantum Machine Learning (QML) models. Inspired by the layered approach of classical neural networks.
As their classical counterpart, Lu et al. [19] highlighted that QNNs are not immune to adversarial attacks. Adversarial attacks involve carefully designed disturbances that can mislead machine learning models, causing them to produce incorrect predictions. Such attacks have long been a significant issue in traditional machine learning [2, 31, 9, 20]. The importance of investigating adversarial attacks on QNNs is reinforced by the no free lunch theorem [24]. This principle suggests that there is not a single model or algorithm universally optimal for all tasks, hinting at inherent vulnerabilities also in quantum-based models. Hence, finding these weak spots is crucial for the resilience and protection of quantum machine learning systems. A further obstacle is that current quantum hardware used to execute these models suffers from noise and decoherence times, which is the reason why Preskill [25] coined the term NISQ-Era for the current state in quantum computing. This means, that due the noise, limited connectivity and decoherence times, inherent in quantum devices, current algorithms can only be executed on smaller toy problems, providing a proof of concept, but suffering from not yet being applicable to practical problems. In this work, we explore the effect of noise, specifically depolarization noise, on the robustness of Quantum Neural Networks (QNNs), aiming to minimize compromises on accuracy and training data. We assumed the structure in Figure 1 as the most realistic scenario, so the input is perturbed on the classical side and then encoded on the quantum circuit. Furthermore, our research builds on the theoretical insights of [6], which suggest that QNNs inherently possess an upper bound of robustness relative to the device’s depolarization noise. This upper bound arises because depolarizing noise, viewed as a stochastic Pauli-channel, effectively scrambles the state of the quantum device, thus blurring information. This blurring leads to a natural tradeoff between accuracy and robustness. We evaluate the effectiveness of this phenomenon by adding multiple depolarizing noise channels to a quantum neural network and comparing its performance with other adversarial training techniques commonly used in classical settings.
[…]
Figure 1: Outline of the process used in this study. Overall, the input is perturbed classically and then encoded on the quantum device using amplitude encoding.
Contrary to the findings of [6], our results indicate that depolarization noise does not improve adversarial robustness in multi-class classifiers. Additionally, we observed that increasing the number of classes diminishes both accuracy and overall robustness, with depolarization noise offering no significant enhancement.
2 Background
2.1 Quantum information and Noise
The fundamental unit of information in quantum computers is the quantum bit (qubit), which is a superposition of the two quantum-mechanical states:
$\psi$ $=$ $\alpha\ket{0}+\beta\ket{1}=\alpha\begin{pmatrix}1\\ 0\end{pmatrix}+\beta\begin{pmatrix}0\\ 1\end{pmatrix},$ (5)
with $~{}{|\alpha|}^{2}+{|\beta|}^{2}=1.$ (6)
Unlike classical computers that use bits with values of either 0 or 1, quantum computers can simultaneously represent both states. This allows for inherent parallel processing in computations. However, when concluding a calculation on a quantum computer, the resulting quantum-mechanical state needs to be measured to convert the calculation result into classical bits. Due to the superposition, the result of the measurement is not definite, but instead the values 0 and 1 of are obtained with a certain probability. The case of Eq. 6, |$\ket{0}$ is measured with a probability of ${|\alpha|}^{2}$ and |$\ket{1}$ with ${|\beta|}^{2}$. Gate-based quantum computers employ gates to perform operations on one or multiple qubits. Gates are unitary matrices. A sequence of multiple gates acting on one or more qubits is called a quantum circuit.
Despite those advantages, current quantum devices suffer from noise. We will now proceed to describe the quantum system and its interaction with the environment. Let us consider that the pure quantum system is described by a density matrix $\rho$. Thus, we can describe the incoherent noise, so the noise from the environment by a linear and completely trace preserving map $\varepsilon(\rho)$. Hereby a convenient strategy is to employ the Kraus operators, so a set of operators ${K_{i}}$, with the property: $\sum_{i}K_{i}^{\dagger}K_{i}\leq I$  [^1] Note that $A^{\dagger}$ describes the conjugate transpose of the matrix $A$.. With those operators, the quantum channel can be described as:
$\varepsilon(\rho)=\sum_{i}K_{i}^{\dagger}\rho K_{i}$ (7)
This concept can describe errors naturally occurring on the quantum device, like Phase-Flips, Bit-Flips or depolarization noise. Since in this work we will focus on the effect of depolarization noise in quantum channels, we will solely describe it here, omitting both phase and bit-flips. We refer the interested reader towards Nielsen and Chuang [23] for further details.
For a single qubit, the Kraus operators $K_{i}$ representing the depolarizing channel can is given in terms of the Pauli matrices $I$, $X$, $Y$, and $Z$, and the depolarizing parameter $p$. The parameter $p$ usually denotes the probability that an error happens. The Kraus operators for the depolarizing channel are:
  • $K_{0}=\sqrt{1-p}I$
  • $K_{1}=\sqrt{\frac{p}{3}}X$
  • $K_{2}=\sqrt{\frac{p}{3}}Y$
  • $K_{3}=\sqrt{\frac{p}{3}}Z$
Here, $I$ is the identity operator and $X$, $Y$, and $Z$ are the standard Pauli matrices:
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\;Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\;Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
The depolarizing noise is a type of quantum noise that affects qubits. It is a model for the process by which a qubit, with some probability, "forgets" its state and becomes a completely mixed state.
2.2 Quantum Neural Network
In the setting of quantum computing, QNNs are realized trought what is known as a variational quantum algorithm (VQA) [3]. The execution of these algorithms is a blend of quantum and classical processes. On the one hand, the learning model is described entirely by a quantum circuit or integrates a quantum circuit at a certain point. The optimization on the other hand is managed using classical methods. A central element in this quantum algorithmic structure is the parametrized quantum circuit (PQC). Composed of unitary transformations, for example including rotation gates:
$R_{z}(\theta)$ $=\begin{pmatrix}e^{-i\theta/2}&0\\ 0&e^{i\theta/2}\\ \end{pmatrix},$ (10)
$R_{y}(\theta)$ $=\begin{pmatrix}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\\ \end{pmatrix},$ (13)
these circuits can be adjusted via the parameters $\theta$ affecting the rotation gates. Regarding training, the loss function is computed using the true labels and the measurements of the quantum device. However, the task of parameter updates to minimize this loss is handled classically, primarily due to its efficient performance in this task.
In our work we will assume a QNN, which consists out of three major parts. At first, we encode the individual datapoints $x\in\mathbb{R}^{n}$ onto the quantum circuit. Hereby a common choice is amplitude encoding using the Mottonen state encoding [22]. Amplitude encoding assumes the normalized input vector and provides a quantum circuit, which will encode the individual elements of $x$ onto the amplitudes of the quantum device:
|$\ket{\psi_{x}}=\frac{1}{||x||}\sum_{i=1}^{n}x_{i}\ket{i}.$ (14)
This strategy is quite attractive, since it provides an exponential reduction of qubits for encoding our datapoints. As a next step a circuit ansatz with the classically optimizable parameters is provided. For this ansatz, numerous options exist [12, 4, 34, 29]. We select the ansatz of Schuld et al. [29] to be architecture-wise in line with Du et al. [6]. The last part of the quantum neural network describes the measurement; the various classes are typically are encoded on different basis states. Here, a set of qubits is selected to be measured. The measured outcomes a assigned to class probabilities. With those probabilities, the loss of the network can be computed and the parameters can be optimized by a classical optimizer. Note that a computation of the gradient is possible on a quantum device, by the so called "parameter-shift" rule.
2.3 Quantum Adversarial Robustness
[…]
Definition 1 (Certified Robustness)
We say $f(x)\in\mathbb{R}^{n}$ is certifiable robust with perturbation budget $\epsilon$ if
$\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x)=\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x+\delta),$ (15)
[…]
In the context of quantum computing, a study by Lu et al. [19] found that QNNs can fall prey to these targeted perturbations. As part of their work, the quantum versions of the Fast Gradient Sign Method (FGSM) [9] and Projected Gradient Descent (PGD) [20] were introduced. Let us denote $x\in\mathbb{R}^{n}$ as (classical) features, $\mathcal{L}$ as loss function incorporating the QNN weights and $\nabla_{x}$ as gradient. In the context of PGD, we define the number of iterations as $T$ and the adversarial budget as $\epsilon$. Formally, we can obtain an adversarial example by iteratively computing:
$x_{adv}^{t+1}=\Pi_{x+\mathcal{S}}(x_{adv}^{t}+\frac{\epsilon}{T}\cdot\mathrm{sign}(\nabla_{x}\mathcal{L})),$
where $\Pi_{x+\mathcal{S}}(x^{\prime})$ clips the value of $x^{\prime}$ towards the space of adversarial perturbations around $x$. After $T$ iterations, the adversarial example can be found as $x_{adv}^{(T)}$.
FGSM on the otherhand, assumes $T=1$ and hence computes once:
$x_{adv}=\Pi_{x+\mathcal{S}}(x+\epsilon\cdot\mathrm{sign}(\nabla_{x}\mathcal{L})),$
Typically the space of adversarial pertubations is defined by a Norm-Ball. This Norm-Ball on which PGD and FGSM attack the network is specified by selecting a specific process for $\Pi_{x+\mathcal{S}}(x^{\prime})$. When using the sign function, the examples will act on the $L_{\infty}$-Norm. The fact that QNNs are vulnerable to adversarial perturbations motivates the discussion about specific measures to counteract adversarial examples. In this work, we will utilize adversarial accuracy as in Def. 2 to describe the network’s vulnerability.
Definition 2 (Adversarial Accuracy)
Let $\mathcal{D}=\{x_{i},{y}_{i}\}_{i=1,\dots,N}$ be a dataset and $a(\cdot,f,\epsilon):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be an adversarial model that tries to find an adversary within the perturbation budget $\epsilon$. Then, the adversarial accuracy of the classifier $f$ on $\mathcal{D}$ under attack $a(\cdot,f,\epsilon)$ is defined as:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}({y}_{i}=f(a(x_{i},f,\epsilon)))$ (16)
Where $\mathbbm{1}$ denotes the indicator function which corresponds to 1 if the argument is true, else is returning 0.
2.4 Adversarial Defense Strategies
Throughout recent years, multiple strategies have been developed to make models more robust against adversarial examples. In this context, we focused on two adversarial defense strategies: adversarial training and randomized smoothing
Adversarial Training
Adversarial training is a technique to improve the robustness of machine learning models against adversarial examples. It involves training the model on a mixture of regular and adversarially perturbed data, allowing the model to learn from and adapt to these perturbations. This method has been extensively studied, with seminal works by Goodfellow et al. [9], Kurakin et al. [15] and Madry et al. [21] being particularly influential in demonstrating its effectiveness and exploring its limits in various contexts.
Randomized Smoothing
Randomized smoothing, not only a method for robustness certification but also an adversarial training technique [27], involves adding noise to training data. This noise, typically Gaussian or Uniform with zero mean, mathematically expressed as $x=x+\delta$, with $\delta\sim\mathcal{N}(0,\sigma^{2}I)$ or $\delta\sim\mathcal{U}(-\sigma,\sigma)$, alters the training data. Unlike standard adversarial training that learns the worst-case decision boundary, randomized smoothing, guides the classifier to learn a smoothed decision boundary. This approach integrates Gaussian perturbations with adversarial examples, not just enhancing empirical robustness but also significantly boosting certifiable robust accuracy of the smoothed classifiers.
2.5 Quantum Certified Robustness
In the context of QML, the robustness bound is motivated by so called quantum differential privacy. Similarly towards classical differential privacy, quantum differential privacy assumes two datapoints $\sigma$ and $\rho$, separated by a small trace-distance: $\tau(\sigma,\rho)=\sqrt{1-\text{Tr}(\sigma,\rho)}$ and for all measurable sets $S$ it must hold that:
$\text{Pr}\left(M(\sigma)\in S\right)\leq e^{\epsilon}\text{Pr}\left(M(\rho)\in S\right)+\delta$ (17)
Where $(\epsilon,\delta)$ is referred to as the privacy budget of the algorithm. $M(x)$ is the quantum classifier in our context. Intuitively, differential privacy in this sense describes that the output of the algorithm differs only slightly due to a small change of input information. In this context, Du et al. [6] have provided a robustness bound for quantum classifiers based on depolarization noise by assuming two quantum states, $\sigma$, and $\rho$, which are the encoded data. The trace-distance, between those two states is then upper-bounded by $\tau_{D}$: $\tau(\sigma,\rho)\leq\tau_{D}$. Note that throughout this work, we assume the infinite sampling case of their upper bound since all models are based on state-vector simulation rather than actual sampling of a quantum simulator or physical device. Next, assume that we are provided with the class probability $y_{C}(\sigma)$ of our class $C$  [^2] Assume $C=0$ without loss of generality., which we can obtain by measuring the final circuit. We define the follow-up class $y_{k\neq C}(\sigma)=\operatorname*{arg\,max}_{k\neq C}\left\{y_{k}(\sigma)\right\}$ and can relate those two classes via quantum differential privacy by Eq. 18.
$y_{C}(\sigma)>e^{2\epsilon}y_{k\neq C}(\sigma)$ (18)
$\epsilon=\ln\left(1+D_{meas}\nicefrac{{(1-p)\tau_{D}}}{{p}}\right)$ (19)
Where $\epsilon$ is defined as Eq. 19. In this term, $p=1-\prod_{i=1}^{j}(1-p_{i})$ is the multiplication of all probabilities of depolarization noise occurrences in each depolarization channel (refer to Figure 2 for an example) and $D_{meas}\geq K$ serves as an upper bound of the dimensionality $K$ of the measurement operator. To clarify this, assume we are provided with $2$ measurements as in Figure 2, our measurement operator would have a dimensionality of $42^{2}=$ 4.
[…]
Figure 2: Example Circuit with deploarization channel $\varepsilon_{p_{1}}$ and two final measurements
3 Robustness Analysis
As mentioned earlier, our goal is to analyze the robustness of QNNs and the effect of the bound provided by Du et al. [6]. Henceforth, we examined the following three research questions as relevant:
  1. 1.
    What is the effect of depolarizing noise on the adversarial accuracy?
  2. 2.
    How does this effect compare to state-of-the-art defense strategies in our setting?
  3. 3.
    Will a blend of depolarizing noise and state-of-the-art defense strategies, have an effect?
To answer those questions, we utilize two datasets: Iris [7], which was motivated from the usage of Du et al. [6], and MNIST [16]. We train our models by the ansatz proposed by Schuld et al. [29] and encode our data on the quantum circuit using amplitude encoding [22].
3.1 Data-Preprocessing
For Iris, we followed the pre-processing of Du et al. [6], which resulted in the removal of the versicolor species and the petal_width feature. For the resulting dataset, we reported 100 feature vectors $x\in\mathbb{R}^{3}$, which were padded by a 00 and normalized by their respective $L_{2}$-Norm. Splitting  [^3] Splitting of the Iris dataset was done at random, utilizing sklearn’s train_test_split function. the dataset into a train and test set resulted in 60 training samples and 40 test samples. As a last step, we one-hot encoded class labels to make them compatible with the resulting measurement outcome from the quantum device.
For the MNIST10 dataset, we rescaled the $28\times 28$ images into smaller images of the size $16\times 16$. Flattening the image resulted in a feature vector $x\in\mathbb{R}^{256}$, which can be realized with eight qubits. Next, we normalized the individual elements of the feature vector to be in the range $[0;1]$ and finally split the data into a training set of $48.000$ images, a validation set of $10.800$ images, and a test set of $1200$ images. To provide additional insight, specific classes of the dataset were subselected to mimic a binary classification task, MNIST2, and a four-class multiclassification task, MNIST4. For MNIST2, we used the classes for the digits one and nine. For MNIST4, we extended the present labels in MNIST2, providing overall the labels one, nine, seven, and three. MNIST2 consisted of $12.691$ datapoints, split  [^4] Splitting of MNIST2, MNIST4 and MNIST10 where done at random, utilizing sklearn’s train_test_split function twice once with with a test_size argument of 0.2 for training and test/validation set and once with a test_size of 0.1 for separating the test and validation set. into $10.152$ training images, $2285$ validation images, and $254$ testing images. Analougus MNIST4, consisting of $25.087$ images, was split into train-, test- and validation-set with the respective sizes $20.069$, $4516$, and $502$.
3.2 Model Architecture
[…]
Figure 3: Strongly Entangling Layer architecture by Schuld et al. [29]. In this quantum circuit outlined for three qubits. Note that the control and target qubits of the entangling layer may vary up to a permutation by different layers.
As pointed out earlier, we used the ansatz of Schuld et al. [29], which consists of strongly entangling layers outlined in Figure 3. Those layers typically consist of three rotation gates, in our case, a combination of $R_{z}$ and $R_{y}$ rotations and a sequence of CNOT-Gates, which are permuted differently on the wires for each layer. The reason for our choice of this particular ansatz is that it provides a simple yet adaptive structure. The number of required qubits and measurements can be determined from the input data size. In contrast, the number of layers influences the amount of parameters and the model’s capacity. One can find the hyper-parameters used in this study in Table 1. The Table also shows the resulting circuit depth, as well as the single- and two-qubit gate count for each circuit, resulting from the selected hyperparameters.
Table 1: Hyperparameters used for the various networks for the different datasets. Note that for the circuit estimate one needs to also include the amplitude encoding procedure by Mottonen et al. [22].
Hyper-Parameter Iris MNIST2 MNIST4 MNIST10 learning rate 0.05 0.005 0.005 0.005 batch size 30 512 512 128 Nr. epochs 100 30 30 30 Nr layers 2 40 40 40 Nr. measurements 1 1 2 4 Nr. Parameters 13 961 961 961 Circuit Depth 7 196 196 196 Single Qubit Gates 12 960 960 960 Two Qubit Gates 4 320 320 320
We want to re-emphasize that specifically for MNIST10, a different architecture for an ansatz specifically exists, which achieved a ${}90\%$ test accuracy [34]. Nevertheless, we wanted to provide a coherent architecture throughout this work to reasonably compare the QNNs among datasets.
Additionally, due to the noisy architecture and limited availability of quantum devices, we have simulated the quantum device using statevector simulation. Specifically, we utilized the auto-grad feature of pytorch and the pennylane library to formulate the ansatz  [^5] Pennylane 0.29.1 and pytorch 2.0.0.
3.3 Robustness Upperbound
During our analysis, we want to compare the bound obtained by Du et al. [6] with empirical results found in our study. Therefore, we utilize Eq. 18 and rewrite it towards obtaining a lower bound for $\epsilon$:
/yC(σ)yC ≠k(σ)$<e^{2\epsilon}$
$\frac{1}{2}\ln\left(\nicefrac{{y_{C}(\sigma)}}{{y_{C\neq k}(\sigma)}}\right)<\epsilon$
Hence, this lower bound for $\epsilon$ is given as follows:
$\epsilon_{min}=\frac{1}{2}\ln\left(\nicefrac{{y_{C}(\sigma)}}{{y_{C\neq k}(\sigma)}}\right)$ (20)
Combining Eq. 20 and Eq. 19, we can obtain the following:
$\epsilon_{min}=\ln\left(1+D_{meas}\nicefrac{{(1-p)\tau_{D}}}{{p}}\right)$ (21)
We can rewrite this to obtain the certified trace distance in terms of the quantum states as follows:
$e^{\epsilon_{min}}=1+D_{meas}\nicefrac{{(1-p)\tau_{D}}}{{p}},$
$\frac{e^{\epsilon_{min}}-1}{D_{meas}}\cdot\frac{p}{(1-p)}=\tau_{D}$
Hence, we can obtain the certified distance $\tau_{D}$ as follows:
$\tau_{D}=\frac{e^{\epsilon_{min}}-1}{D_{meas}}\cdot\frac{p}{(1-p)}$ (22)
Assuming that the probability of depolarization noise $p$ is known, we can see that the certified distance $\tau_{D}$ is only dependent on the fraction of the class probability and to the follow-up class, the size of the measurement operator.
In our study, we will attack the quantum classifier with the FGSM and PGD attacks [^6] FGSM for the MNIST-Classifiers and PGD for the Iris Classifier. We utilized the FGSM attack since computing the gradient of the quantum circuit concerning the input is quite resource-intensive when a noise channel is evaluated., both bounded by the $L_{\infty}$-norm. Hence, we need to ensure that the certified distance we will obtain from our classification is an upper bound to those two norms. Du et al. [6] have already argued, that the $L_{2}$-Norm is bounded by the trace distance:
$L_{2}\leq\tau(\sigma,\rho)$ (23)
Consequently, we can utilize $L_{\infty}\leq L_{2}$ to argue that this bound is also valid for attacks using an $L_{\infty}$-Norm. We did not normalize the data points in the MNIST datasets by the $L_{2}$-norm but performed normalization during the encoding. A comparison of this bound towards the classical epsilon value is hence difficult, since attacks will be made in the unnormalized space. Nevertheless, the upper bound can provide us with an estimate of the certifiable robustness of the data points and a certain degree of explanatory power. To make this more clear, assume a correct classification, which will yield the relation: $y_{C}(\sigma)>y_{C\neq k}(\sigma)$. We can conclude from this relation, that the fraction $\nicefrac{{y_{C}(\sigma)}}{{y_{C\neq k}(\sigma)}}$ will be greater than one: $\nicefrac{{y_{C}(\sigma)}}{{y_{C\neq k}(\sigma)}}>1$. Hence, the natural logarithm in Eq. 21 will also be greater or equal to zero. Since $e^{0}=1$, approaching $e^{\epsilon_{min}}-1$ will result in a positive upper bound. On the other hand, in case we are provided with a misclassification, the class probability will be smaller than the follow-up class: $y_{C}(\sigma)<y_{C\neq k}(\sigma)$. Hence, the fraction $\nicefrac{{y_{C}(\sigma)}}{{y_{C\neq k}(\sigma)}}$ will be smaller than one, and the natural logarithm will be negative. This will result in $e^{\epsilon_{min}}-1<0$; so, the certifiable distance will be smaller than 00. We interpret this as it is impossible to certify the data point based on the model’s output.
Note that this reasoning indicates that a model with more parameters and good generalization can be more robust. This may be achieved by quantum convolutional neural networks or a different design of a variational ansatz.
4 Experiments
With our experiments, we try to answer the three questions raised above. At first, we want to see the influence of depolarization noise on a multi-class classifier. Secondly, our interest shifts towards whether one can defend our QNNs with other adversarial training methods, specifically randomized smoothing and adversarial training using PGD. Finally, we want to elaborate on whether we can chain adversarial training with depolarization noise as a blend of two adversarial defense mechanisms.
Table 2: Test accuracy ($Acc_{test}$) for the different models trained during this study. For the iris dataset we reused the test as validation set during training due to the limited available datapoints. Additionally the mean certified distance is reported for the test set, with respect to different depolarization noise values: $[0.1,0.5,0.9]$. We marked the best mean certified distance per experiment in bold and the negative certified distances as red.
Dataset Train Param. Acc. Certified Distance: $\tau_{D}$ (×$\times 10^{-2}$)
$\epsilon$ or $\sigma$ $p=0.1$ $p=0.5$ $p=0.9$
Iris vanilla 100.0 4.176 5.960 1.092
PGD 0.01 100.0 2.823 5.215 1.291
0.05 100.0 2.811 5.184 1.276
0.09 100.0 2.774 5.100 1.254
RS 0.1 100.0 2.824 5.181 1.273
0.5 97.50 2.661 4.412 1.029
0.9 70.00 2.375 3.502 0.723
MNIST2 vanilla 99.61 2.429 0.150 0.006
PGD 0.01 99.61 2.366 0.135 0.003
0.05 99.61 2.372 0.127 0.001
0.09 99.21 2.392 0.123 0.002
RS 0.1 99.61 2.369 0.124 0.001
0.5 99.61 2.378 0.116 -0.003
0.9 99.61 2.253 0.111 0.002
MNIST4 vanilla 90.44 1.035 0.042 -0.007
PGD 0.01 90.64 0.997 0.034 -0.006
0.05 90.64 0.995 0.044 -0.003
0.09 90.24 1.000 0.037 -0.005
RS 0.1 91.24 0.988 0.048 -0.004
0.5 91.04 0.988 0.047 -0.003
0.9 91.43 0.916 0.050 -0.003
MNIST10 vanilla 80.00 0.205 -0.012 -0.007
PGD 0.01 80.50 0.172 -0.049 -0.014
0.05 80.08 0.173 -0.049 -0.015
0.09 81.75 0.164 -0.030 -0.011
RS 0.1 79.50 0.180 -0.062 -0.017
0.5 76.17 0.174 -0.053 -0.014
0.9 71.91 0.167 -0.013 -0.006
4.1 Training of the Models
Throughout this study, we trained four types of QNNs: one for the Iris dataset, one for the MNIST-2, one for the MNIST-4, and one for the MNIST-10 dataset. We train all models using Adam [14] as an optimizer by varying the batch size, learning rate, and number of epochs. We will refer to those models as vanilla models. We used categorical-cross entropy implementation from pytroch as a loss function for Iris, MNIST2 and MNIST4. For MNIST10, we found that simply using negative log likelihood ($l(y,\hat{y})=-\sum_{i}y_{i}\log(\hat{y}_{i})$) as a loss further increased the accuracy, which is why this loss function was used. Concerning adversarial training, we train the models using the PGD attack with $L_{\infty}$-norm for the Iris dataset with $T=50$ steps. Due to performance issues, we train the MNIST-2, MNIST-4, and MNIST-10 models with $T=3$ steps. We additionally employ randomized smoothing during training; therefore, we perturb the original image with noise from a uniform distribution and use this as a training sample. We report the variance of this training as $\sigma$, and the mean is set to zero. We employ early stopping for both randomized smoothing and adversarial training. Therefore, we use the loss value, at which we considered the models to be converged. For Iris this value was set to $0.6$. For MNIST2, we set this value to $0.45$, for MNIST4 to $1.15$ and for MNIST10 to $1.7$. Table 2 reports the test accuracy, combined with the mean certifiable distance $\tau_{D}$ we derived from the test-set.
4.2 Effect of Depolarization Noise on a Multi-Class Classifier
We compare the vanilla networks of the Iris, MNIST2, MNIST4, and MNIST10 datasets for depolarization noises $p$ in the range $[0.1,0.5,0.9]$. We want to point out that we have included a depolarizing noise of $0.3$ for MNIST2. Our results are shown in Figure 4.
Overall, the two binary classification networks, MNIST2 and Iris, perform well in terms of plain accuracy on the test-set. Depolarization noise might help here in the sense that it improves the classifier for larger epsilon values, which we can observe for $p=0.5$ and for $p=0.9$ for the Iris dataset. Similar behavior can also be seen in Du et al. [6, Apdx. H], where the adversarial accuracy was evaluated under a $L_{2}$-Norm. For MNIST2, we did not find any significant improvement in terms of adversarial accuracy. Similarly, for MNIST4 and MNIST10, we can observe that depolarization noise will not increase the robustness. We account this fact towards two things.
The first issue concerns the limited capacity of the network to learn the data distribution. This observation is supported by Table 2. We can see that the accuracy further decreases with an increase in the number of classes. Additionally, when looking at the certified bounds, we can first see that the minimum of all computed bounds is typically negative, and the maximum trace distance is typically smaller than for the binary classification. The negative bound can be attributed to our inability to achieve a $100\%$ accuracy on the test set.
(a) Iris
(b) MNIST2
(c) MNIST4
(d) MNIST10
Figure 4: Robustness accuracy for the various vanilla classifiers for Iris, MNIST2, MNIST4 and MNIST10.
Secondly, we have measured two qubits to obtain the classes for MNIST4 and four qubits for MNIST10. Another attribution of the bad performance could be, that depolarization noise will this way automatically increase the probability of the follow-up class since we are assigning two classes to one qubit. Nevertheless, this assumption requires a further experimental and analytical evaluation, which is subject for future work.
4.3 Effect of Adversarial Training
Our results, regarding the adversarial training of the QNNs can be found in Figure 5. Overall, we can see, that neither of the techniques provides a reasonable estimate of enhancing the robustness of the QNN. Additionally, we can see a similar behavior as shown for the depolarizing noise, specifically that the networks tend to get less robust for larger classes. This development shows a parallel trend towards the fact that the testing accuracy drops around $\approx 10\%$ from MNIST2 to MNIST4 and around $\approx 10\%$ from MNIST4 to MNIST10. Interestingly, randomized smoothing and adversarial training tend not to worsen the robustness accuracy.
(a) Iris: PGD
(b) Iris: RS
(c) MNIST2: PGD
(d) MNIST2: RS
(e) MNIST4: PGD
(f) MNIST4: RS
(g) MNIST10: PGD
(h) MNIST10: RS
Figure 5: Comparsion of the two methods randomized smoothing (RS) and adversarial training (PGD) for a set of parameters on the QNN. Adversarial training was conducted with PGD($L_{\infty}$, $T=50$) for the Iris classifier, PGD($L_{\infty}$, $T=3$) for all MNIST classifiers
4.4 Effect of Depolarization Noise to Adversarial Training
Since adversarial training and the effect of depolarization noise tend to behave similarly, we want to explore if adding depolarization noise on top of an adversarial training procedure can improve the overall robustness of the network. Therefore, we look closer at the MNIST2 dataset and enhance the adversarially trained QNN circuits with depolarization noise while attacking them. Our results can be observed in Figure 6. Overall, we find that adding depolarization noise on top of adversarial training does not show the ability to provide improvements for the MNIST2 Dataset. We can see a similar picture for other datasets when looking at the certified trace distances on the test set in Table 2.
(a) MNIST2: Adversarial Training
(b) MNIST2: Randomized Smoothing
Figure 6: Comparison of the adversarially trained QNNs with respect to different depolarization noise.
5 Related Work
Since quantum neural networks are prune to adversarial attacks [18, 19], over recent years, multiple authors have tried to use quantum computing techniques to enhance adversarial robustness for both classical and quantum neural networks. Consequently, Sahdev and Kumar [26] have utilized a quantum phase estimation to provide a quadratic speedup for randomized smoothing. Nevertheless, since quantum phase estimation is not feasible to be executed on NISQ-Hardware and they additionally required a nested quantum phase estimation for specifically quantum neural networks, we refrained from utilizing this method as a comparison for the studies in our work. Gong et al. [8] proposed a vanishing gradient for introducing adversarial noise on a quantum circuit. Nevertheless, the method of Gong et al. [8] assumes that the data is in a quantum state, which is then adversarially perturbed during the quantum circuit. We focused on encoding classical data on the quantum circuit rather than assuming already encoded quantum states. Hence, in our scenario, the method is hence not applicable. Therefore, we refrained from including Gong et al. [8] in our study. Additionally, Weber et al. [35] have provided a fundamental link between quantum hypothesis testing and binary classification. Their work also shows the connection between classical randomized smoothing and depolarization noise on a quantum circuit. Finally, our approach to train and test the robustness of a multi-class quantum neural network is novel in that the community primarily focuses on binary classification. The only work focusing on the development of variational ansätze for multi-class classifiers is Wang et al. [34], which created a structure to classify MNIST10 with an accuracy of $\approx 90\%$.
6 Conclusion
Evaluating the robustness of Quantum Neural Networks (QNNs), we addressed three primary research questions.
Firstly, we assessed the impact of depolarization noise on a multi-class classifier. It was observed that introducing more classes to the QNN decreases both training and test accuracy. Furthermore, the overall robustness of the QNN diminishes and is not improved by depolarization noise. This trend is particularly evident when examining the certifiable distances $\tau_{D}$; for all MNIST datasets, $\tau_{D}$ decreases with an increase in depolarizing noise and even tends towards negative values when $p=0.9$ for MNIST4 and MNIST10. Secondly, exploring alternative methods for adversarial training of the network also failed to yield benefits, as did the addition of depolarization noise to adversarially trained networks.
It is important to note that Du et al. [6] asserted that the certifiable distance depends solely on the overall depolarizing noise, the dimensions, and the ratio of the true class probability to the probability of the subsequent class. Our work demonstrates that increasing depolarizing noise in a quantum channel negatively impacts both this certifiable distance and the adversarial accuracy in practice. Consequently, future research should focus on obtaining weights that enhance class separation in the QNN context. This is crucial because neither the number of measurements nor the depolarizing noise in a quantum device can be controlled. This premise initially motivated our exploration of certifiable distance. Hereby it would be interesting to lean on the work of Liao et al. [17], to find more suitable architectures than the one used in this work.
Aside from enhancing quantum neural networks to achieve reasonable performance, a secondary objective could be to replicate this study with other widely-used datasets, such as CIFAR-10 or CIFAR-100. This would be of particular interest, since it could allow a quantum-classical comparison of the robustness of QNNs, towards networks present in Croce et al. [5]  [^7] https://robustbench.github.io/.
Finally, we wish to highlight to the reader that, despite these challenges, the foundational assumption of quantum differential privacy and the feasibility of establishing a robustness bound with minimal depolarization noise in the quantum circuit enable the creation of inherently certifiable networks. However, based on our findings, we acknowledge the limited practical applicability of quantum neural networks in their current form.
Acknowledgement
The research is supported by the Bavarian Ministry of Economic Affairs, Regional Development and Energy with funds from the Hightech Agenda Bayern.